OpenStudy (anonymous):

find the area of the following polar equ. r=3(1+cosX) thanks in advancee

OpenStudy (anonymous):

Use the formula: \[\int\limits_{a}^{b}(1/2)r^2d\] But to find a and b, set 0 = 3 + 3cosx and find two values that makes this work.

OpenStudy (anonymous):

\[\int\limits\limits_{a}^{b}(1/2)r^2dr\]

OpenStudy (anonymous):

Make sure to fan me ^^

OpenStudy (anonymous):

already did lol thanks again

OpenStudy (anonymous):

oh yeah it's you hahaha. No problem!